Jefferson Lab HALL A Logbook for March 09 - All EXCEPT Automatic Entries

  • Main INDEX

    Tuesday

  • 266567 03/31/09 11:38 Moffit Obtained 5 LC 1461N from Hall C
  • 266566 03/31/09 11:05 R. Michaels scalers added to R-HRS, problems.
  • Monday

    
    

    Saturday

    
    

    Friday

  • 266558 03/27/09 16:53 Moffit/Mihov Re: Bad DE fanout.
  • 266557 03/27/09 16:44 Moffit/Mihov Bad linear fanout for DE:13-16
  • Wednesday

  • 266556 03/25/09 19:24 V. Sulkosky BigBite Weldment Cables Disconnected on L-HRS
  • 266555 03/25/09 19:19 V. Sulkosky L-HRS Delay Removed for S1 and S2m TDC Signals
  • 266550 03/25/09 12:29 R. Michaels BB scaler (more)
  • 266549 03/25/09 12:25 R. Michaels mods to BB scalers
  • Tuesday

  • 266548 03/24/09 20:05 Moffit Problem found with T3 delay at BB weldment.
  • 266547 03/24/09 17:42 Moffit Scaler changes in BB Trigger Supervisor
  • 266546 03/24/09 16:36 R. Michaels baseline DAQ configs ready
  • 266542 03/24/09 15:50 V. Sulkosky Re: L-HRS Databases Updated for S1, S2m and RF Timing Signals
  • 266541 03/24/09 15:44 V. Sulkosky L-HRS Databases Updated for S1, S2m and RF Timing Signals
  • 266540 03/24/09 15:13 A. Camsonne Compton electron detector chamber opening - Electron detector test
  • 266539 03/24/09 15:12 R. Michaels Rtmp working (HAPPEX config)
  • 266538 03/24/09 13:26 V. Sulkosky RF Signals Moved to 1877 for L-HRS
  • 266537 03/24/09 13:19 V. Sulkosky L-HRS S1 and S2m TDC Change
  • 266536 03/24/09 12:14 Moffit Re: LV Translator
  • 266535 03/24/09 12:02 Moffit BB level translator for TS signals
  • 266534 03/24/09 09:41 R. Michaels HAPPEX DAQ Roc23 in repair, use Rtmp
  • Monday

  • 266533 03/23/09 17:33 A. Camsonne Compton Electron comics run 16462
  • 266532 03/23/09 17:29 A. Camsonne Compton Electron comics run 16461
  • 266531 03/23/09 12:28 Yi Qiang Fake spin flip signal ON: 20s
  • 266530 03/23/09 12:19 Yi Qiang Water calibration during the weekend
  • 266529 03/23/09 08:59 A. Camsonne Fan tray for helicity
  • Saturday

  • 266528 03/21/09 13:48 R. Michaels HAPPEX crate fan failure
  • Friday

  • 266527 03/20/09 18:32 Moffit BB TS crate now on hareboot26
  • 266526 03/20/09 18:31 Moffit Changed scalers in ROC5
  • Wednesday

  • 266519 03/18/09 11:45 Moffit setup ip for CAMAC controller and scope
  • Tuesday

  • 266518 03/17/09 16:55 atrig Trigger-Download
  • 266517 03/17/09 10:41 Moffit Disabled mysql database scripts.
  • 266516 03/17/09 10:41 R. Michaels 1 kHz HAPPEX DAQ testing
  • Monday

  • 266515 03/16/09 17:09 brads Valve on He3 bottle closed
  • 266514 03/16/09 16:57 brads All BB HVs turned off
  • 266513 03/16/09 16:57 R. Michaels planned helicity test tomorrow
  • 266512 03/16/09 16:48 brads BB MWDC HVs off, gas flow shut off
  • 266511 03/16/09 16:48 brads BigBite magnet turned off
  • 266510 03/16/09 15:40 Moffit CAMAC stuff from Hall C
  • 266509 03/16/09 14:06 Yi Qiang All long fibers locked
  • 266508 03/16/09 13:59 brads Low deadtime in run 2212 OK
  • 266507 03/16/09 13:53 brads BB Cerenkov gain measurement with LED
  • 266505 03/16/09 13:48 fansler Post Beam Check list
  • 266503 03/16/09 13:45 Yi Qiang NMR Cross Calibration (late entry)
  • 266500 03/16/09 13:42 D. Parno Run 2212 this morning had low deadtime and L1A glitch
  • 266499 03/16/09 13:39 Yi Qiang Internal Temperature Test of Transverse Pumping (late entry)
  • 266486 03/16/09 12:41 YI Qiang Cooling jets shut down
  • 266474 03/16/09 12:14 D. Parno HAPPEX/Compton runs stopped
  • 266470 03/16/09 11:34 Yi/JP 1 atm N2 in reference cell
  • 266469 03/16/09 11:27 Yi/Jian-ping Reference Cell vented
  • 266468 03/16/09 11:27 Yi/Jian-ping Reference Cell vented
  • 266465 03/16/09 10:49 brads Hall is in Restricted (late entry)
  • 266464 03/16/09 10:48 Yi Qiang EPR with 270 degrees holding field
  • 266463 03/16/09 10:21 Yi Qiang EPR with 90 degrees holding field
  • 266462 03/16/09 09:56 Yi Qiang Holding field adjusted for Negative BigBite polarity
  • 266460 03/16/09 09:50 brads BigBite ramped back up in negative polarity
  • 266458 03/16/09 08:56 brads BB HVs updated to negative polarity settings
  • 266457 03/16/09 08:44 brads BigBite magnet ramped down for switch -> Negative pol.
  • 266455 03/16/09 08:11 miham Restricted access authorization
  • 266454 03/16/09 08:06 miham OWL shift summary
  • 266448 03/16/09 07:05 hyao target to empty
  • 266441 03/16/09 06:16 miham Spot++
  • 266440 03/16/09 05:54 ellie Golden Run set to 2211 (BB)
  • 266427 03/16/09 04:50 ellie Set Golden Runs to 20718 (LHRS) and 2210 (BB)
  • 266426 03/16/09 04:40 hyao Run 2210 Data Check
  • 266425 03/16/09 04:19 miham Negative deadtime
  • 266418 03/16/09 04:08 miham HRS-L momentum change
  • 266411 03/16/09 02:25 hyao Restart CODA for LHRS
  • 266410 03/16/09 02:24 hyao NMR=58.7%
  • 266402 03/16/09 02:12 ellie VDC Efficiencies Screen
  • 266397 03/16/09 01:14 ellie Set Golden Runs to 20711 (LHRS) and 2203 (BB)
  • 266391 03/16/09 00:05 LeRose Shift Summary (swing)
  • 266390 03/16/09 00:04 LeRose Shift Summary (swing)
  • Sunday

  • 266386 03/15/09 23:59 LeRose Left Arm to 1.19 GeV/c
  • 266381 03/15/09 22:35 ellie Gold Run set to 20709 (LHRS) and 2199 (BB)
  • 266378 03/15/09 22:32 LeRose BigBite HV trip
  • 266376 03/15/09 21:53 brads BB HVs uploaded for positive polarity (late entry)
  • 266375 03/15/09 21:49 brads BB Cosmics 2197 will be hard to interpret
  • 266362 03/15/09 21:01 LeRose Beam Back
  • 266361 03/15/09 20:55 Moffit F1TDC reference plots
  • 266360 03/15/09 20:47 brads BB Thresholds set for production
  • 266359 03/15/09 20:46 Moffit Access: What I did.. F1TDC reference channels in Slots 6,8,9
  • 266357 03/15/09 20:43 Yi Qiang Holding field adjusted for positive BigBite polarity
  • 266354 03/15/09 20:40 LeRose BigBite changed from negative to positive polarity
  • 266353 03/15/09 20:34 brads BigBite in positive polarity, back to 710A
  • 266352 03/15/09 20:27 LeRose Left Arm set to 1.34 GeV/c
  • 266351 03/15/09 20:23 LeRose Left Q2/3 cycled
  • 266349 03/15/09 20:08 brads BigBite ramped down for polarity change.
  • 266348 03/15/09 20:04 xqian NMR (60.96%)
  • 266341 03/15/09 18:36 Kalyan BigBite shower high voltages need to be changed after BB is switched to positive polarity @8pm
  • 266340 03/15/09 18:30 LeRose Left Trip
  • 266339 03/15/09 18:21 LeRose Left Dipole Back
  • 266337 03/15/09 17:58 LeRose Left Dipole Trip
  • 266333 03/15/09 17:48 D. Parno Syncroc glitch - Run 2179
  • 266328 03/15/09 16:26 brads Quit and restarted the software -- EPICS problem fixed
  • 266327 03/15/09 16:19 brads EPICS gateway problem?
  • 266325 03/15/09 16:04 Yi Qiang Day Shift Summary
  • 266323 03/15/09 16:00 Yi Qiang BigBite HV tripped
  • 266321 03/15/09 15:50 D. Parno Low deadtime for Run 2190
  • 266320 03/15/09 15:47 D. Parno Syncroc glitch an artifact of ROOT plot?
  • 266319 03/15/09 15:47 Yi Qiang 6x5.2 (MCC value) raster since Mar 14, 18:02
  • 266318 03/15/09 15:43 D. Parno Syncroc glitch in Run 2190
  • 266317 03/15/09 15:41 Yi Qiang Raster size increase to 6x5.2 MCC value
  • 266316 03/15/09 15:38 D. Parno More L1A doubling glitches in Run 2190
  • 266313 03/15/09 15:35 Kalyan NMR --> Polarization 61.24%
  • 266310 03/15/09 15:10 Yi Qiang 1/3 decreasing of Raster in X direction
  • 266309 03/15/09 14:50 D. Parno Sample LHRS VDC Efficiency
  • 266308 03/15/09 14:50 Yi Qiang BPM chart
  • 266307 03/15/09 14:49 Yi Qiang Raster size check done, ramp beam to 15uA
  • 266304 03/15/09 14:44 D. Parno L1A glitch in Run 2185
  • 266303 03/15/09 14:43 Yi Qiang Raster IOC fixed, but No BPM readings
  • 266302 03/15/09 14:31 D. Parno LHRS golden run -> 20701
  • 266301 03/15/09 14:24 Yi Qiang iochla2 for Raster is Down
  • 266300 03/15/09 14:14 no_name MCC IOC problem
  • 266297 03/15/09 13:56 Yi Qiang Hall A Tools
  • 266296 03/15/09 13:55 Yi Qiang Spot++ of run 20704
  • 266283 03/15/09 12:55 A. Camsonne Restarted Compton
  • 266278 03/15/09 11:39 Yi Qiang Target->3He
  • 266275 03/15/09 11:14 Yi Qiang N2 cell in beam pressure: 116 psig
  • 266272 03/15/09 11:02 Yi Qiang Target->N2 for reference run
  • 266267 03/15/09 10:07 Yi Qiang Set to Kin 4F
  • 266266 03/15/09 10:06 Kalyan NMR --> Polarization 60.65%
  • 266259 03/15/09 08:44 Kalyan BB DAQ problem solved
  • 266257 03/15/09 08:34 Kalyan BB DAQ problem
  • 266256 03/15/09 08:29 Kalyan end-run failed -- dead time 100%
  • 266254 03/15/09 08:03 miham OWL shift summary
  • 266248 03/15/09 06:48 R. Michaels Target polarization vs time
  • 266243 03/15/09 06:16 miham Spot++
  • 266242 03/15/09 06:12 miham End-of-run file missing
  • 266241 03/15/09 06:07 miham Sync Check problem
  • 266240 03/15/09 06:05 miham BB Cer ADCs shifted
  • 266237 03/15/09 05:35 R. Michaels NMR (60.24)
  • 266226 03/15/09 05:14 R. Michaels guess about DT bug
  • 266225 03/15/09 05:04 R. Michaels neg. deadtime appears benign
  • 266220 03/15/09 04:15 R. Michaels scaler check in run 2176
  • 266216 03/15/09 04:05 R. Michaels negative deadtime recurs
  • 266212 03/15/09 03:40 R. Michaels data quality ok
  • 266211 03/15/09 03:16 D. Parno T8 problems began halfway through Run 2163
  • 266208 03/15/09 02:58 R. Michaels spot++
  • 266203 03/15/09 02:47 D. Parno Big Bite shower threshold returned to -110 mV for beam
  • 266202 03/15/09 02:39 R. Michaels T8 the culprit of all problems
  • 266197 03/15/09 02:18 D. Parno Oscillations in L1A signal for Run 2167
  • 266193 03/15/09 02:16 D. Parno BB Shower threshold set to -20 mV
  • 266192 03/15/09 02:06 R. Michaels request access
  • 266190 03/15/09 02:05 R. Michaels NMR (57.95)
  • 266188 03/15/09 02:00 D. Parno L1A for Run 2164
  • 266187 03/15/09 01:58 D. Parno F1 Reference Signals and syncroc gap in Run 2164
  • 266184 03/15/09 01:58 R. Michaels infamous test.dat problem
  • 266183 03/15/09 01:55 D. Parno More Run 2164 Problems
  • 266182 03/15/09 01:51 R. Michaels tried power cycling TS12 -- no help
  • 266177 03/15/09 01:30 miham Negative deadtime in online analysis
  • 266176 03/15/09 01:19 miham Missing peaks in BB Cerenkov data
  • 266173 03/15/09 01:11 R. Michaels L1A double-pulsing
  • 266172 03/15/09 00:56 R. Michaels negative deadtime
  • 266167 03/15/09 00:38 D. Parno Timing Signals from Crate 5 Slot 11 Look OK
  • 266166 03/15/09 00:18 D. Parno Signals were from Crate 5 Slot 8, not Crate 8
  • Saturday

  • 266160 03/14/09 23:53 D. Parno BB Timing Signals from Crate 8 Look Okay
  • 266159 03/14/09 23:48 hyde Evening Shift Summary
  • 266158 03/14/09 23:47 hyde HRS-L replay
  • 266155 03/14/09 23:21 Chiranjb NMR @ 11:15 P
  • 266152 03/14/09 22:53 Chiranjib target graphs
  • 266151 03/14/09 22:47 D. Parno F1 reference channel problem goes away in cosmics run
  • 266148 03/14/09 21:45 D. Parno BB Shower Sum TDC Signal Changes with F1 Reference Structure
  • 266147 03/14/09 21:43 hyde BB DC trip
  • 266145 03/14/09 21:36 hyde beam /BB HV trip
  • 266135 03/14/09 21:22 Chiranjb Laser Spectrum
  • 266134 03/14/09 21:21 Chiranjib NMR @ 9:15 P
  • 266129 03/14/09 21:15 hyde BCA
  • 266128 03/14/09 20:43 brads F1 ref issue NOT fixed by reboot...
  • 266126 03/14/09 20:28 brads Power cycled ROC5
  • 266124 03/14/09 20:23 brad F1 reference channel problem for slot 8, 11; runs 2107+
  • 266121 03/14/09 20:06 adaq Hall A tools
  • 266120 03/14/09 20:00 A. Camsonne Compton Edet run 16408
  • 266119 03/14/09 19:52 posik golden->20687, 2156
  • 266118 03/14/09 19:50 A. Camsonne Compton Edet run 16406 16407
  • 266117 03/14/09 19:48 hyde HRS-L replay
  • 266116 03/14/09 19:47 posik F1 clock chek low
  • 266113 03/14/09 19:28 A. Camsonne Compton Edet run 16405
  • 266112 03/14/09 19:24 A. Camsonne Compton Edet run 16404
  • 266111 03/14/09 19:22 Yi/Chiranjib NMR before production @ 1.34 Gev/c
  • 266110 03/14/09 19:08 hyde harp
  • 266107 03/14/09 18:54 hyde Harp
  • 266104 03/14/09 18:31 hyde BBDC
  • 266101 03/14/09 18:16 hyde Production Raster
  • 266100 03/14/09 18:12 Chiranjib N2 pressure during run 2154 / 20685
  • 266097 03/14/09 18:05 hyde raster study
  • 266096 03/14/09 17:56 Chiranjib Target moved to N2
  • 266095 03/14/09 17:55 Chiranjib SPOT ++
  • 266090 03/14/09 17:48 hyde Half Wave Plate
  • 266089 03/14/09 17:45 Chiranjib Target moved to Be0
  • 266088 03/14/09 17:31 Yi Qiang Air Flow/Heater Power adjustment
  • 266087 03/14/09 17:29 Yi /Chiranji Target ready
  • 266086 03/14/09 17:26 hyde power permit
  • 266085 03/14/09 16:31 Yi/Chiranjib Oven heaters back on-line
  • 266083 03/14/09 15:34 brads Target heater has been replaced
  • 266082 03/14/09 15:33 brads LHRS set to kin 4H (1.34 GeV/c)
  • 266077 03/14/09 15:19 gjin target moved to pickup coils position
  • 266076 03/14/09 13:30 Yi Qiang Cutting tubes....
  • 266067 03/14/09 11:41 Yi Qiang Replacing heater... Estimated time is about 2 hours
  • 266066 03/14/09 11:34 A. Camsonne compton iochacp2 informations
  • 266065 03/14/09 10:49 A. Camsonne RICH CODA TI module borrowed
  • 266064 03/14/09 10:31 A. Camsonne Compton electron dtector spacer removed
  • 266063 03/14/09 10:04 Yi Qiang One Heater down
  • 266062 03/14/09 09:44 Yi Qiang Hall Fully surveyed, oven issue solved
  • 266061 03/14/09 08:47 gjin problem w/ sync
  • 266060 03/14/09 08:30 Yi Qiang 3He Oven Base Heater Down
  • 266059 03/14/09 08:28 flay Double peaks in LHRS Gas Cerenkov TDCs
  • 266053 03/14/09 08:09 miham Owl shift summary
  • 266045 03/14/09 06:38 miham Spot++
  • 266044 03/14/09 06:36 miham Raster size to small
  • 266031 03/14/09 05:17 hyao NMR 62.47%
  • 266026 03/14/09 04:25 miham BB wire chamber trip
  • 266017 03/14/09 02:59 D. Parno LHRS Golden Run -> 20675
  • 266016 03/14/09 02:44 miham Missing End-of-run file
  • 266015 03/14/09 02:32 miham N2 cell pressure
  • 266012 03/14/09 02:23 hyao target to He3
  • 266009 03/14/09 01:55 miham HRS-L momentum change
  • 266006 03/14/09 01:43 hyao target to N2
  • 265999 03/14/09 00:07 gjin target polarization and target oven temperature
  • 265998 03/14/09 00:06 gjin holding current and target cell temperature
  • 265997 03/14/09 00:03 gjin ref cell and target readbacks
  • 265996 03/14/09 00:02 gjin abs_spectr
  • Friday

  • 265994 03/13/09 23:55 D. Parno NMR measurement was during CODA runs
  • 265993 03/13/09 23:49 hyde Eve 13 Mar 09 shift summary
  • 265992 03/13/09 23:44 gjin NMR(62.36%)
  • 265991 03/13/09 23:36 hyde BCA = 0!?
  • 265986 03/13/09 23:13 D. Parno LHRS Golden Run -> 20671
  • 265985 03/13/09 23:06 D. Parno Syncroc Problems: What to Look for
  • 265984 03/13/09 22:53 brads re daq (ts12 problem)
  • 265983 03/13/09 22:49 brads re: DAQ
  • 265982 03/13/09 22:44 D. Parno BigBite DAQ running again
  • 265981 03/13/09 22:43 hyde switch to Kin 4C
  • 265978 03/13/09 22:37 gjin Target Moved to 3He Production
  • 265976 03/13/09 22:29 hyde daq
  • 265975 03/13/09 22:26 hyde DAQ
  • 265974 03/13/09 22:12 gjin change kinematics into 4C
  • 265973 03/13/09 22:06 D. Parno Sync roc detachment is an artifact
  • 265971 03/13/09 22:00 g Target moved to reference cell
  • 265969 03/13/09 21:45 hyde BB synch check
  • 265968 03/13/09 21:34 D. Parno LHRS VDC Efficiency Capture
  • 265955 03/13/09 20:32 R. Michaels clarification of tgt move (Late)
  • 265950 03/13/09 19:44 gjin NMR (62.83%)
  • 265945 03/13/09 19:24 hyde BCA
  • 265944 03/13/09 19:09 posik L1A and sync check run 2137
  • 265931 03/13/09 17:39 hyde HV trip
  • 265930 03/13/09 17:23 A. Camsonne Compton Edet run 16367
  • 265929 03/13/09 17:21 A. Camsonne Compton Edet run 16366
  • 265928 03/13/09 17:17 A. Camsonne Compton Edet run 16365
  • 265923 03/13/09 16:19 D. Parno Compton Polarization History 3/12 - 3/13
  • 265922 03/13/09 16:16 A. Camsonne Compton electron detector runs
  • 265921 03/13/09 16:15 hyde raster spot++
  • 265919 03/13/09 15:59 gjin Target moved to 3He
  • 265914 03/13/09 15:50 Nilanga Shift Summary part 2
  • 265913 03/13/09 15:33 D. Parno Compton cavity mirror tune
  • 265910 03/13/09 15:22 R. Michaels move to N2 ref. cell
  • 265907 03/13/09 14:34 A. Camsonne Compton Edet at 26000
  • 265906 03/13/09 14:24 posik LHRS golden->20659 (kin 4E270)
  • 265903 03/13/09 14:21 brads Energy lock was off (on now)
  • 265900 03/13/09 14:12 A. Camsonne Compton Edet dark current
  • 265889 03/13/09 13:20 Nilanga Reat check with large raster
  • 265888 03/13/09 13:11 LeRose Target to 3He
  • 265887 03/13/09 13:08 Nilanga spot++
  • 265885 03/13/09 13:07 Nilanga Shift Summary
  • 265883 03/13/09 12:50 R. Michaels NMR (63.99)
  • 265882 03/13/09 12:31 R. Michaels beam back, put in BeO
  • 265881 03/13/09 11:46 flay Compton End of Run Failed
  • 265880 03/13/09 11:30 brads Moller run summary for d2n
  • 265879 03/13/09 10:56 flay T1 vs. T2 trigger for runs 1665, 2107 (EDIT)
  • 265878 03/13/09 10:55 flay T1 vs. T2 trigger for runs 1665, 2107
  • 265877 03/13/09 10:37 flay Re: Deadtime too high?
  • 265876 03/13/09 10:36 R. Michaels RF zone done (no beam)
  • 265875 03/13/09 10:32 flay LHRS Gas Cerenkov TDCs -- Double Peaks?
  • 265874 03/13/09 10:23 Nilanga Deadtime too high ?
  • 265871 03/13/09 10:06 R. Michaels runs with diff. raster
  • 265864 03/13/09 09:38 R. Michaels raster check 2x2.
  • 265863 03/13/09 09:36 Nilanga Compton HV trip
  • 265862 03/13/09 09:33 R. Michaels rates not const. vs raster size
  • 265857 03/13/09 09:28 Nilanga Raster check with 3He
  • 265848 03/13/09 09:12 Nilanga spot++
  • 265845 03/13/09 09:12 R. Michaels moving to He3
  • 265842 03/13/09 08:57 brads Restarted Compton/Happex DAQs
  • 265841 03/13/09 08:47 R. Michaels ask for 2 uA
  • 265840 03/13/09 08:10 R. Michaels access for polarity change
  • 265837 03/13/09 08:00 Chiranjib SHIFT SUMMARY
  • 265829 03/13/09 06:43 Chiranjib/ H ROC 3 , ROC 4 , TS11 rebooted
  • 265821 03/13/09 05:51 Chiranjib TARGET ----->> He-3 @ P_0 = 0.8GeV/c
  • 265816 03/13/09 05:35 Chiranjib Target moved to N2
  • 265804 03/13/09 04:13 A. Camsonne Compton electron detector Edet run
  • 265799 03/13/09 03:24 Chiranjib TARGET ----->> He-3 @ P_0 = 1.34 GeV/c
  • 265788 03/13/09 02:13 Chiranjib N2 run
  • 265783 03/13/09 02:04 Chiranjib Target moved to N2
  • 265782 03/13/09 02:03 Chiranjib NMR @ 2 A
  • 265779 03/13/09 01:58 posik vertex reconstruction
  • 265778 03/13/09 01:54 Chiranjib Laser Absorption Spectrum
  • 265777 03/13/09 01:40 A. Camsonne Compton Edet test run 16340
  • 265774 03/13/09 01:36 A. Camsonne Compton Edet test run 16338
  • 265772 03/13/09 01:35 A. Camsonne Compton Edet test run 16337
  • 265766 03/13/09 01:20 Chiranjib Hall A Tools
  • 265759 03/13/09 00:25 flay Vertex Reconstruction
  • 265754 03/13/09 00:15 posik LHRS double adc peak
  • Thursday

  • 265750 03/12/09 23:58 xqian Swing Shift Summary
  • 265749 03/12/09 23:45 hyde HRS
  • 265744 03/12/09 23:35 posik golden run LHRS->20621 BB->2107
  • 265732 03/12/09 23:07 brads Shoulder in BB vertex reconstruction histo?
  • 265731 03/12/09 23:04 Kalyan N2 pressure increased to 116psig after a while in beam
  • 265727 03/12/09 22:51 Kalyan Reference cell to N2 - pressure 115 psig (in beam)
  • 265726 03/12/09 22:51 Yi Qiang The flucation is normal
  • 265725 03/12/09 22:39 A. Camsonne Shielding effect on Compton Edet
  • 265722 03/12/09 22:33 brads Cerenkov PMTs 9, 10 same as before amplifier change
  • 265721 03/12/09 22:24 Kalyan Target screens
  • 265716 03/12/09 22:03 posik BB run 2017 L1A check
  • 265715 03/12/09 21:36 Kalyan Reference cell pressure - 117psig
  • 265714 03/12/09 21:32 hyde Target error
  • 265703 03/12/09 21:18 hyde Raster, Trigger rates
  • 265702 03/12/09 21:14 Kalyan NMR (polarization = 65.34%)
  • 265701 03/12/09 21:13 D. Parno Compton rates and spectrum are good
  • 265698 03/12/09 20:59 brads re: Harp scan looks OK
  • 265697 03/12/09 20:50 Kalyan Laser 3 bits flutuating (?)
  • 265696 03/12/09 20:46 xqian HARP SCAN results
  • 265695 03/12/09 20:43 Kalyan Target to optics
  • 265694 03/12/09 20:05 O. Glamazdin Moller results
  • 265693 03/12/09 19:51 Kalyan Laser Absorption Spectrum
  • 265692 03/12/09 18:56 A. Camsonne Compton cavity tune vertical setting
  • 265691 03/12/09 18:55 A. Camsonne Compton cavity mirror position
  • 265690 03/12/09 18:10 D. Parno Compton HV Trip
  • 265689 03/12/09 18:02 ellie We have beam!
  • 265688 03/12/09 17:50 Ole Hansen Day Shift Summary
  • 265687 03/12/09 17:15 Jack Target flammable gas sensor
  • 265681 03/12/09 15:54 gjin Target moved to empty position
  • 265680 03/12/09 13:52 gjin NMR
  • 265671 03/12/09 13:05 brads re: Yellow light on the Target Flammable Gas Sensors
  • 265670 03/12/09 10:40 A. Camsonne Compton rate problem Hall C current variable
  • 265669 03/12/09 10:23 brads Experiment update
  • 265668 03/12/09 09:11 ellie LHRS at P0=1.12GeV/c
  • 265666 03/12/09 08:08 miham Owl Shift Summary
  • 265664 03/12/09 08:01 gjin NMR
  • 265655 03/12/09 04:03 hyao NMR (61.21%)
  • 265652 03/12/09 00:53 miham Yellow light on the Target Flammable Gas Sensors
  • 265651 03/12/09 00:03 Solvignon Shift summary
  • Wednesday

  • 265649 03/11/09 22:58 Solvignon NMR Measurement: 51.7% (no beam).
  • 265648 03/11/09 22:51 brads re: LHRS peaks on top of each other
  • 265639 03/11/09 21:18 Solvignon Laser Absorption Spectrum
  • 265638 03/11/09 21:13 brads BB Thresholds set to production values
  • 265635 03/11/09 21:08 brads BB Trigger restored to Sh.AND.PSh
  • 265634 03/11/09 21:00 brads adaql2 rebooted, io scheduler changed to 'deadline'
  • 265628 03/11/09 20:47 ellie BB Cer TDC Screen Grab (Run 2102)
  • 265627 03/11/09 20:29 Yi Qiang Target Polarization keep dropping: incorrect spin state
  • 265624 03/11/09 20:13 ellie LHRS Replay Peaks on Top Of Each Other
  • 265623 03/11/09 20:03 Yi Qiang Target Spin Rotated to 270 degrees, Field: Transverse +
  • 265622 03/11/09 19:55 Yi Qiang More Masing test
  • 265617 03/11/09 19:18 Yi Qiang EPR Calibration with Transverse Pumping
  • 265615 03/11/09 18:32 brads There was a power glitch in Cebaf Center around 5pm
  • 265614 03/11/09 17:38 Kalyan Run 2102 is a cosmics after HV update
  • 265612 03/11/09 17:35 Kalyan BigBite HV updated for negative polarity run
  • 265611 03/11/09 17:29 Yi Qiang Interlock glitch caussed laser trip
  • 265610 03/11/09 17:28 R. Michaels cabling and DAQ changes.
  • 265609 03/11/09 17:26 R. Michaels HAPPEX DAQ deadtime checks
  • 265608 03/11/09 16:23 flay LHRS PRL coefficients
  • 265606 03/11/09 16:01 gilman day shift summary 11 march 2009
  • 265605 03/11/09 14:25 brads BB thresholds set to -20mV for Cosmics
  • 265604 03/11/09 14:23 brads PSh/Sh logic set to OR for T6 Cosmics <-- MUST be restored before 9pm
  • 265603 03/11/09 14:09 Yi Qiang After field adjustment: NMR: 60.98%
  • 265602 03/11/09 13:59 Yi Qiang Field adjusted for BigBite Negative polarity
  • 265593 03/11/09 12:40 LeRose Absorption Spectrum
  • 265592 03/11/09 11:58 A. Camsonne Compton vacuum after ion pump turned on
  • 265591 03/11/09 11:22 riordan BigBite Polarity to Negative
  • 265590 03/11/09 10:47 riordan Status of Machine - No beam until Thursday
  • 265589 03/11/09 10:05 gjin MWDC check
  • 265588 03/11/09 09:37 LeRose NMR (57.10%)
  • 265587 03/11/09 08:27 D. Parno Owl shift summary 3/11
  • 265576 03/11/09 05:13 D. Parno Absorption spectrum
  • 265575 03/11/09 04:31 Parno/Frankl NMR Measurement (57.1%)
  • 265573 03/11/09 00:11 gilman swing shift summary Mar 10, 2009
  • Tuesday

  • 265572 03/10/09 23:06 brads Compton vacuum snapshot
  • 265571 03/10/09 21:55 brads Compton line pump-down update . . . beamline valves closed until tomorrow?
  • 265562 03/10/09 21:21 riordan Summary for March 9 Swing - March 10 Day
  • 265561 03/10/09 20:14 Moffit Access Activity
  • 265560 03/10/09 17:45 brads Pump down of Compton beamline is /real/ slow.
  • 265559 03/10/09 16:34 Parno/Frankl FADC Compton Asymmetries
  • 265558 03/10/09 16:03 Saha Shift Summary
  • 265556 03/10/09 15:52 Solvignon NMR measurement: 57.12% (no beam)
  • 265555 03/10/09 15:28 brads re: LHRS Gas Cerenkov
  • 265554 03/10/09 14:50 flay LHRS Gas Cerenkov
  • 265553 03/10/09 14:50 flay LHRS Gas Cerenkov
  • 265544 03/10/09 12:42 Solvignon Laser Absorption Spectrum
  • 265543 03/10/09 11:24 gjin MWDC hit efficiency
  • 265542 03/10/09 11:17 gjin MWDC check
  • 265541 03/10/09 09:12 riordan Experiment Status
  • 265540 03/10/09 08:19 gilman March 10 owl shift summary
  • 265538 03/10/09 06:40 W. Troth Laser Absorption Spectrum Owl Shift
  • 265537 03/10/09 06:38 W. Troth NMR Measurement (56.73%)
  • 265528 03/10/09 01:33 brads BB Trigger Sh.AND.PSh; T1, T6 Thresholds set to -110mV (BB ready for production)
  • 265526 03/10/09 01:05 W. Troth Temporary Error with NMR Measurement
  • 265525 03/10/09 00:57 W. Troth NMR Measurement (56.85%)
  • 265524 03/10/09 00:43 brads Run 2101 is cosmics calibration for final BB positive-polarity HV setting
  • 265523 03/10/09 00:33 Kalyan Re: BigBite HV change for positive polarity runs
  • Monday

  • 265519 03/09/09 23:17 brads Run 2100 is BB Cosmics with T1 trigger only (positive polarity)
  • 265516 03/09/09 23:04 brads Added 'F1 Clock Check' page to detectorBB.cfg
  • 265515 03/09/09 22:33 brads Cosmics run afte 1st pass HV gain match for BigBite (run 2099)
  • 265514 03/09/09 22:32 brads Updated halogscript to 1.45 (URL cleanup)
  • 265512 03/09/09 22:18 Kalyan BigBite HV changed for positive polarity runs
  • 265511 03/09/09 22:02 brads T6 trigger set to PSh.OR.Sh for cosmics run after HV change.
  • 265510 03/09/09 22:00 brads New version of the online_C.so (onlana GUI)
  • 265501 03/09/09 21:13 A. Camsonne Dark current after board modification for electron detector
  • 265500 03/09/09 20:42 adaq Cosmic run 2097 with T1
  • 265499 03/09/09 20:38 Kalyan T6 trigger back to Sh.AND.PSh
  • 265488 03/09/09 19:51 A. Camsonne Injector BCM calibration done
  • 265485 03/09/09 19:28 A. Camsonne I0L02 vs faraday cup logfile
  • 265481 03/09/09 19:15 A. Camsonne I0L02 vs faraday cup calibration
  • 265480 03/09/09 18:13 riordan Summary for March 8 Swing - March 9 Day
  • 265479 03/09/09 17:14 brads Cosmics run with T6 trigger
  • 265478 03/09/09 17:10 brads LHRS @ 0.8 GeV, BigBite ready to go
  • 265477 03/09/09 17:02 brads PSh/Sh logic set to OR for T6 Cosmics <-- MUST be restored before 9pm
  • 265472 03/09/09 16:36 A. Camsonne Compton electron detector bias voltage calibration
  • 265471 03/09/09 16:23 A. Camsonne Compton electron detector swapped distribution board and nano ampere meter.
  • 265470 03/09/09 16:12 A. Camsonne Compton electron detector threshold DAC calibration check
  • 265467 03/09/09 15:52 Yi Qiang Field adjusted for BigBite positive polarity
  • 265464 03/09/09 14:59 gilman 9 March 2009 day shift summary
  • 265463 03/09/09 14:24 Moffit copies of F1 TDC reference channels..
  • 265462 03/09/09 14:17 R. Michaels daylight-savings halog bug-fix
  • 265451 03/09/09 12:57 Yi Qiang Abort Masing test on 3He Target due to BB polarity change
  • 265450 03/09/09 12:50 flay BB 2088 Cosmic Run
  • 265449 03/09/09 12:29 gilman current status
  • 265448 03/09/09 12:28 gjin MWDC check
  • 265447 03/09/09 12:27 gjin MWDC hit efficiency
  • 265446 03/09/09 12:11 gilman HRSL set to +1.6 GeV/c
  • 265445 03/09/09 11:43 Jack (Techs) BigBite power supply water leak
  • 265444 03/09/09 11:43 gilman no beam until ?
  • 265440 03/09/09 11:28 riordan BigBite T1 threshold set to -20 mV
  • 265439 03/09/09 11:13 brads BB Cerenkov PMT 9, 10 are bypassing the 1st amplifier
  • 265438 03/09/09 10:41 riordan In Controlled Access for BigBite and LHRS polarity change
  • 265433 03/09/09 09:16 jpchen NMR (63%)
  • 265431 03/09/09 09:12 jpchen
  • 265430 03/09/09 09:09 gilman raster size check results
  • 265427 03/09/09 08:59 jpchen
  • 265424 03/09/09 08:54 gilman spot++ check
  • 265422 03/09/09 08:47 gilman HARP scan results
  • 265421 03/09/09 08:46 jpchen
  • 265420 03/09/09 08:46 brads Restarted Happex/Compton(x2) DAQs
  • 265419 03/09/09 08:31 flay LHRS Golden Run set to 20607
  • 265417 03/09/09 08:03 zhangyi owl shift summary
  • 265416 03/09/09 07:53 zhangyi halla main tools
  • 265415 03/09/09 06:20 W. Troth Target moved to Empty Target
  • 265414 03/09/09 06:19 W. Troth Target moved to Empty Target
  • 265413 03/09/09 05:54 W. Troth Laser Absorption Spectrum Owl Shift 05:52
  • 265404 03/09/09 03:09 W. Troth NMR Measurement (60.63%)
  • 265399 03/09/09 02:06 W. Troth NMR Measurement (61.3%)
  • 265394 03/09/09 00:58 brads halogscript bug
  • 265393 03/09/09 00:45 brads Sunday's Halog entries will return at 1am...
  • 265391 03/09/09 00:04 bogdanw shift summary
  • 265390 03/09/09 00:01 brads LHe cryo control fixed (we hope!)
  • Sunday

  • 265389 03/08/09 23:59 riordan More work done regarding cryo levels
  • 265386 03/08/09 23:53 LeRose Target changed to 3He
  • 265383 03/08/09 23:33 LeRose Ref Cell RTD's
  • 265380 03/08/09 23:02 brads Dipole LHe PID screwed up again?
  • 265379 03/08/09 22:01 LeRose Laser Absorption Spectrum
  • 265378 03/08/09 21:58 LeRose Target Reference Cell
  • 265377 03/08/09 21:54 bogdanw HRS dipole trip
  • 265376 03/08/09 21:50 LeRose NMR (60.47%)
  • 265373 03/08/09 21:35 xqian L1A for 2079 and 2080
  • 265364 03/08/09 21:19 bogdanw BHWP is "OUT"
  • 265359 03/08/09 21:07 S. Sirca Sync check (run 2080)
  • 265356 03/08/09 20:27 D. Parno Compton DAQ restored
  • 265354 03/08/09 20:17 riordan 1L04 Vacuum really rising
  • 265353 03/08/09 20:16 flay BB 2078 snyc roc and L1A check
  • 265350 03/08/09 20:08 bogdanw LHRS dipole trips again
  • 265348 03/08/09 19:15 S. Sirca General tools screen
  • 265347 03/08/09 19:14 S. Sirca spot++ check
  • 265344 03/08/09 18:50 D. Parno Compton polarization history 3/6 - 3/8
  • 265343 03/08/09 18:40 D. Parno Messy L1A for Run 2077
  • 265340 03/08/09 18:04 riordan Summary for March 7 Swing - March 8 Day
  • 265339 03/08/09 17:22 LeRose Cryo Problems Resolved (We Hope!)
  • 265337 03/08/09 17:06 LeRose NMR (60.99%)
  • 265336 03/08/09 17:01 riordan Left Q1 and Dipole Cryo Levels since 9:30
  • 265331 03/08/09 16:55 LeRose Left Dipole Back
  • 265330 03/08/09 16:39 LeRose Left Dipole Trip
  • 265327 03/08/09 16:22 posik forgot last plot BB cer analysis
  • 265326 03/08/09 16:21 flay Re: Cer Analysis
  • 265325 03/08/09 16:18 posik cer analysis for BB runs 2019,2020 and 2024
  • 265324 03/08/09 16:09 posik cer analysis for BB runs 2019,2020 and 2024
  • 265321 03/08/09 15:56 X. Jiang day shift summary
  • 265318 03/08/09 15:34 X. Jiang Left arm dipole tripped again.
  • 265316 03/08/09 15:25 flay Snyc Roc and L1A Check for BB 2072
  • 265315 03/08/09 15:07 riordan Re: [Tech] Cryo fluctuations on LHRS magnets
  • 265311 03/08/09 14:59 X. Jiang left diploe reset done.
  • 265310 03/08/09 14:34 X. Jiang no_key
  • 265308 03/08/09 14:31 flay Compton Trouble
  • 265307 03/08/09 14:13 flay L1A vs. Clock Count for BB 2071
  • 265304 03/08/09 13:45 brads [Tech] Cryo fluctuations on LHRS magnets
  • 265303 03/08/09 13:30 X. Jiang left arm Q1 He level low,
  • 265302 03/08/09 13:19 flay LHRS E/P Calibration
  • 265293 03/08/09 13:13 riordan 1L04 Vacuum continues to rise
  • 265290 03/08/09 13:01 X. Jiang BB HV tripped, while beam had a trip.
  • 265285 03/08/09 12:32 posik aux daqs restarted (compton trouble)
  • 265282 03/08/09 11:45 Holmstrom NMR Measurement (61.36%)
  • 265281 03/08/09 11:44 posik BB 2068 sync gap/dead time/L1A
  • 265278 03/08/09 11:14 X. Jiang spot++ for run 20596
  • 265277 03/08/09 11:05 gjin MWDC efficiency check
  • 265273 03/08/09 10:06 X. Jiang started production run again
  • 265271 03/08/09 10:03 Holmstrom 3He target position
  • 265268 03/08/09 09:48 Holmstrom Moved target to optics
  • 265267 03/08/09 09:24 X. Jiang IOC problem still not fixed.
  • 265266 03/08/09 08:49 X. Jiang controlled access to reboot IOC-HLA.
  • 265265 03/08/09 08:47 X. Jiang Update on IOC
  • 265264 03/08/09 08:46 posik Dead time for runs 2059-2065
  • 265262 03/08/09 08:04 Solvignon Shift summary
  • 265261 03/08/09 07:46 Solvignon IOC down: no beam
  • 265254 03/08/09 05:34 W. Troth NMR Measurement: 59.77% with beam on
  • 265237 03/08/09 04:15 W. Troth Laser Absorption Spectrum Owl Shift 03/08/2009 04:14
  • 265234 03/08/09 03:39 D. Parno Syncroc check -- Run 2061
  • 265233 03/08/09 03:38 Solvignon Spot++ (checklist)
  • 265232 03/08/09 03:34 hyao sync check 2 for run 2061
  • 265231 03/08/09 03:07 W. Troth NMR Measurement: 58.65% with beam on (13.3uA)
  • 265229 03/08/09 03:04 Solvignon Beam back
  • 265227 03/08/09 01:26 Solvignon No beam for an hour
  • 265223 03/08/09 00:28 D. Parno L-arm VDC efficiency capture
  • 265222 03/08/09 00:23 Solvignon BB HV trip
  • 265218 03/08/09 00:08 D. Parno Deadtime/L1A for Run 2058 (14 uA)
  • 265216 03/08/09 00:05 bogdanw shift sumamry
  • Saturday

  • 265213 03/07/09 23:58 A. Rakhman NMR (56.93%)
  • 265212 03/07/09 23:57 D. Parno Re: BB HV trip during 2058
  • 265209 03/07/09 23:52 D. Parno LHRS good electron rate for G kinematic
  • 265208 03/07/09 23:48 ycwang BigBite HV tripped during 2058
  • 265205 03/07/09 23:03 riordan DT and L1A for 14 uA
  • 265204 03/07/09 23:02 D. Parno Higher deadtime and messy L1A at 14 uA
  • 265203 03/07/09 22:47 D. Parno LHRS good electron rate for 4I kinematic
  • 265198 03/07/09 22:13 riordan DT and L1A rate at 13uA
  • 265197 03/07/09 21:46 ycwang Sync Check1(straight lines)
  • 265196 03/07/09 21:39 D. Parno LHRS golden run set to 20588
  • 265195 03/07/09 21:22 Yi Qiang NMR Before Production: 51.34%
  • 265192 03/07/09 21:17 A. Rakhman Target is in He3 position
  • 265189 03/07/09 21:10 ycwang Sync Check(straight lines)
  • 265188 03/07/09 20:46 bogdanw kin#4g tools
  • 265185 03/07/09 20:43 A. Rakhman target is in reference cell
  • 265184 03/07/09 20:37 ycwang BeO is good
  • 265183 03/07/09 20:36 Yi Qiang Polarization after rotation is ~ 47%
  • 265182 03/07/09 20:33 A. Rakhman Target position is BeO
  • 265181 03/07/09 20:23 Yi Qiang Target Spin Rotated to 90 degrees: Transverse -
  • 265180 03/07/09 20:22 A. Rakhman Target laser status
  • 265171 03/07/09 20:17 Yi Qiang COMET laser power test
  • 265170 03/07/09 20:05 ycwang When increased raster from 5.2X6.5 to 6X7, the rates increase approx. 50%
  • 265163 03/07/09 19:26 brads Crosshairs on BeO are washed out by laser in longitudinal config
  • 265162 03/07/09 19:12 ycwang Sync Check1(Straight lines)
  • 265161 03/07/09 18:59 A. Rakhman Laser Absorption Spectrum
  • 265160 03/07/09 18:52 bogdanw raster in last run for kin #4I
  • 265159 03/07/09 18:43 riordan BeO Crosshairs Not in Viewer
  • 265154 03/07/09 17:46 D. Parno Compton Polarization History 3/6-3/7
  • 265153 03/07/09 17:01 A. Rakhman Target Position
  • 265152 03/07/09 16:38 riordan Summary for March 6 Swing - March 7 Day
  • 265151 03/07/09 16:16 A. Rakhman Target screen snapshots
  • 265150 03/07/09 16:12 A. Rakhman Laser Absorption Spectrum swing shift 3/7/2009
  • 265149 03/07/09 16:10 A. Rakhman NMR (57.46%)
  • 265146 03/07/09 16:05 Solvignon Shift summary
  • 265142 03/07/09 15:58 Solvignon Beam off for 1/2 hour
  • 265139 03/07/09 15:44 flay LHRS Cerenkov Pedestals Changed
  • 265136 03/07/09 15:25 posik L1A check for BB run 2050
  • 265135 03/07/09 15:16 Nuruzzaman Screenshot of Sync check1 for run# 2050
  • 265132 03/07/09 14:31 Nuruzzaman Screenshot of Sync check1 for run# 2049
  • 265129 03/07/09 14:22 posik L1A fo BB run 2048
  • 265128 03/07/09 14:09 posik L1A check for BB run 2046
  • 265125 03/07/09 13:34 Nuruzzaman Screenshot of Sync check1 for run# 2048
  • 265124 03/07/09 13:21 flay Deadtime for run# 2046
  • 265105 03/07/09 12:10 Nuruzzaman Screenshot of Sync check1 for run# 2045
  • 265104 03/07/09 12:00 A. Deur NMR (57.32%)
  • 265103 03/07/09 11:54 A. Deur Lasers in remote control now
  • 265101 03/07/09 11:53 Nuruzzaman Screenshot of Sync check1 for run# 2044
  • 265095 03/07/09 11:07 Nuruzzaman Screenshot of Sync check1 for run# 2043
  • 265092 03/07/09 10:03 Nuruzzaman Screenshot of Sync check1 for run# 2042
  • 265089 03/07/09 09:17 Nuruzzaman Screenshot of Sync check1
  • 265088 03/07/09 08:57 flay Run# 2040 MWDC
  • 265086 03/07/09 08:52 flay Deadtime
  • 265084 03/07/09 08:30 A. Deur Laser Current in local and readout is zero
  • 265083 03/07/09 08:24 Nuruzzaman Screenshot of Sync check1
  • 265082 03/07/09 08:13 A. Deur Absorption Spectrum: ref is not good?
  • 265080 03/07/09 08:03 X. Jiang owl shift summary
  • 265077 03/07/09 08:01 riordan Changed Beam Half Wave Plate to IN
  • 265076 03/07/09 08:00 W. Troth NMR Measurement: 57.45% with beam on (8uA)
  • 265073 03/07/09 07:39 miham Sync. gap in run #2039
  • 265070 03/07/09 07:20 X. Jiang End of Run 20575
  • 265067 03/07/09 07:03 miham Wrong end-of-run entry
  • 265066 03/07/09 06:57 miham Sync. gap in run #2038
  • 265063 03/07/09 06:22 miham SPOT++
  • 265061 03/07/09 05:56 miham Sync. gaps in run #2037
  • 265060 03/07/09 05:50 W. Troth Laser Absorption Spectrum Owl Shift 3/7/2009
  • 265056 03/07/09 05:10 miham Sync. gaps in run #2036
  • 265053 03/07/09 04:29 miham No sync. gaps in run #2035
  • 265042 03/07/09 03:25 miham Sync. Gaps in run #2034
  • 265039 03/07/09 02:39 miham ER1 warning
  • 265038 03/07/09 02:33 D. Parno Syncroc check - Run 2033
  • 265037 03/07/09 02:21 W. Troth NMR Measurement: 58.44% with beam on (15.9uA)
  • 265032 03/07/09 02:01 D. Parno Syncroc check - Run 2032
  • 265029 03/07/09 01:14 D. Parno Syncroc check - Run 2029
  • 265026 03/07/09 00:59 miham BigBite wire chambers
  • 265025 03/07/09 00:56 D. Parno Odd Structure in LHRS VDC TDC
  • 265022 03/07/09 00:47 Parno/Camson Compton DAQ is Back
  • 265021 03/07/09 00:28 X. Jiang chamber trips and beam trips are correlated ???
  • 265020 03/07/09 00:25 adev BigBite WireChamber Measured Voltage change!
  • 0 03/07/09 00:23 no_name no_key
  • 265017 03/07/09 00:20 miham BB wire chamber trip
  • 265015 03/07/09 00:15 adev BigBite WireChamber Measured Voltage change!
  • 265013 03/07/09 00:00 D. Parno Syncroc check, Run 2028
  • Friday

  • 265012 03/06/09 23:59 S. Sirca Swing shift SUMMARY
  • 265011 03/06/09 23:58 D. Parno Compton DAQ trouble
  • 265010 03/06/09 23:41 Holmstrom Scalers
  • 265007 03/06/09 23:10 riordan Summary for March 4 swing - March 5 Day
  • 265006 03/06/09 22:41 adev BigBite WireChamber Measured Voltage change!
  • 265004 03/06/09 22:30 adev BigBite WireChamber Measured Voltage change!
  • 265003 03/06/09 22:27 D. Parno Syncroc check - Run 2026
  • 265000 03/06/09 22:19 adev BigBite WireChamber Measured Voltage change!
  • 264998 03/06/09 22:09 adev BigBite WireChamber Measured Voltage change!
  • 264997 03/06/09 22:05 D. Parno Big Bite golden run set to 2025
  • 264994 03/06/09 21:37 D. Parno Syncroc check: Run 2025
  • 264993 03/06/09 21:16 S. Sirca Hall A General Tools after beam back
  • 264992 03/06/09 21:14 Solvignon Laser Absorption Spectrum: much better absorption in Pol.3He position
  • 264989 03/06/09 21:10 Solvignon NMR Measurement: 60.88% with beam on (15uA)
  • 264988 03/06/09 21:08 S. Sirca spot++ check with He3, raster 5.2x6.5
  • 264985 03/06/09 20:38 Solvignon Target in 3he position
  • 264983 03/06/09 20:35 adev BigBite WireChamber Measured Voltage change!
  • 264982 03/06/09 20:34 S. Sirca spot++ check with BeO, raster 2.5x3.2
  • 264980 03/06/09 20:30 D. Parno Compton laser tune
  • 264979 03/06/09 20:26 O. Glamazdin Moller measurements result
  • 264978 03/06/09 20:20 Solvignon Laser Absorption Spectrum
  • 264969 03/06/09 20:09 Solvignon Target in BeO+Optics position
  • 264968 03/06/09 19:31 A. Camsonne Low current Edet run 16240
  • 264967 03/06/09 19:25 flay Change of Pedestals for the LHRS PRL1&2
  • 264966 03/06/09 19:20 A. Camsonne Compton electron detector run 16239
  • 264965 03/06/09 19:11 A. Camsonne Compton electron detector run 16237
  • 264964 03/06/09 19:04 A. Camsonne Compton electron detector run 16236
  • 264963 03/06/09 19:03 A. Camsonne Compton electron detector run 16235
  • 264962 03/06/09 19:03 A. Camsonne Compton electron detector run 16234
  • 264961 03/06/09 19:03 A. Camsonne Compton electron detector run 16233
  • 264960 03/06/09 19:02 A. Camsonne Compton electron detector run 16233
  • 264959 03/06/09 19:01 A. Camsonne Compton electron detector run 16232
  • 264958 03/06/09 18:55 A. Camsonne Compton electron detector run 16231
  • 264957 03/06/09 18:54 brads F1 ref clock check added to online gui
  • 264956 03/06/09 18:50 A. Camsonne Compton electron detector run 16230
  • 264955 03/06/09 18:49 A. Camsonne Compton electron detector run 16229
  • 264954 03/06/09 18:40 A. Camsonne Compton electron detector run 16228
  • 264953 03/06/09 18:37 A. Camsonne Compton electron detector run 16227
  • 264952 03/06/09 18:30 A. Camsonne Low current Edet run 16226
  • 264951 03/06/09 18:30 A. Camsonne Low current Edet run 16225
  • 264950 03/06/09 18:25 A. Camsonne Compton electron detector run 16225
  • 264949 03/06/09 18:02 A. Camsonne Dark current Edet Compton 1 uA
  • 264948 03/06/09 18:01 A. Camsonne Low current Edet run 16225
  • 264946 03/06/09 17:18 D. Parno Compton Asymmetry at 4-Pass Today
  • 264945 03/06/09 17:12 Solvignon Target in empty position
  • 264944 03/06/09 17:00 posik sync gaps in BB run 2024
  • 264943 03/06/09 16:35 adev BigBite WireChamber Measured Voltage change!
  • 264942 03/06/09 16:31 S. Sirca Stop DAQ in order to do the Moller measurement
  • 264939 03/06/09 16:24 posik LHRS golden run kin 4I0->20570
  • 264937 03/06/09 15:57 Ole Hansen Day Shift Summary
  • 264934 03/06/09 15:37 Ole Hansen Start 3He production
  • 264933 03/06/09 15:31 LeRose Target changed to 3He
  • 264930 03/06/09 15:24 posik Comptons and HAPPEX restarted
  • 264929 03/06/09 15:21 LeRose Ref Cell RTD Readbacks
  • 264927 03/06/09 14:51 adev BigBite WireChamber Measured Voltage change!
  • 264926 03/06/09 14:45 Ole Hansen MMWDC HV was off - turning it on
  • 264924 03/06/09 14:44 Nuruzzaman The Target is reference cell
  • 264923 03/06/09 14:43 Nuruzzaman The Target is reference cell
  • 264920 03/06/09 14:23 LeRose Target changed to Reference Cell (with N2)
  • 264919 03/06/09 14:18 Ole Hansen Target to ref, raster 4x4
  • 264918 03/06/09 14:17 Ole Hansen Raster scan
  • 264917 03/06/09 14:09 LeRose Abs_spectr_030609_Day(2)
  • 264916 03/06/09 14:04 LeRose NMR Sweep Shot , polarization 62.86%
  • 264915 03/06/09 14:01 LeRose Target to 3He
  • 264913 03/06/09 14:00 Ole Hansen spot++ run 20568
  • 264911 03/06/09 13:54 Ole Hansen CODA problems?
  • 264908 03/06/09 13:47 Yi Qiang First NMR after rotated to Longitudinal: ~ 64%
  • 264907 03/06/09 13:46 Ole Hansen spot check
  • 264906 03/06/09 13:41 Target Group Spin direction rotated to 0 degree
  • 264905 03/06/09 13:38 Ole Hansen Harp results
  • 264903 03/06/09 13:23 Ole Hansen Beam ready
  • 264902 03/06/09 13:06 adev BigBite WireChamber Measured Voltage change!
  • 264901 03/06/09 12:56 Ole Hansen BB MWDC HV off
  • 264900 03/06/09 12:56 LeRose Target changed to Optics
  • 264899 03/06/09 12:17 Ole Hansen LHRS to 1.42 GeV/c
  • 264898 03/06/09 12:17 Ole Hansen Beam off for pass change
  • 264883 03/06/09 11:51 D. Parno Compton Polarization Measurements During Unpolarized 4-Pass Running
  • 264880 03/06/09 11:35 LeRose Abs_spectr_030609_Day
  • 264879 03/06/09 11:32 D. Parno Compton orbit lock restored
  • 264878 03/06/09 11:30 LeRose NMR Sweep Shot , polarization 60.87%
  • 264873 03/06/09 11:02 brads Killing runaway hvcli process on adaql1
  • 264872 03/06/09 11:01 adev BigBite WireChamber Chamber Status change!
  • 264871 03/06/09 11:00 D. Parno Compton orbit lock was not on
  • 264870 03/06/09 10:52 D. Parno Drift in Compton chicane vertical position
  • 264867 03/06/09 10:26 gjin routine MWDC check
  • 264866 03/06/09 10:24 gjin MWDC hit efficiency
  • 264863 03/06/09 09:14 riordan TDC Spectra Look Good
  • 264861 03/06/09 09:05 riordan Power Cycled ROC5
  • 264859 03/06/09 08:59 Nuruzzaman Error in replay for BB run #2016
  • 264858 03/06/09 08:57 riordan Scintillator and sum time spectrum now flat
  • 264857 03/06/09 08:47 Moffit Sum and Right Scint in same F1TDC.
  • 264855 03/06/09 08:40 riordan Rebooted BigBite ROCs
  • 264854 03/06/09 08:36 brads BB Tscint - Tsum double peaks
  • 264850 03/06/09 08:31 Ole Hansen Beam back
  • 264849 03/06/09 08:27 brads Double peaks in 10 Row-Sum TDCs too.
  • 264848 03/06/09 08:20 Ole Hansen Beam gone
  • 264844 03/06/09 08:01 Gao Owl shift summary
  • 264841 03/06/09 07:25 riordan New Structure in BigBite Right Scint TDCs
  • 264840 03/06/09 07:24 miham Problem with BB scint. TDC
  • 264839 03/06/09 07:19 riordan Sync Signals don't match in 2009 briefly
  • 264838 03/06/09 07:05 Solvignon NMR measurement: 61.86% (beam on)
  • 264837 03/06/09 07:04 Solvignon NMR measurement: 63.17% (beam on)
  • 264821 03/06/09 04:41 Solvignon Spot++: 3x4 raster
  • 264820 03/06/09 04:30 Solvignon Laser Absorption Spectrum
  • 264809 03/06/09 03:20 miham Rate tests
  • 264804 03/06/09 03:04 Solvignon Spot++ for 4x4
  • 264801 03/06/09 02:53 Solvignon Target in 3He position
  • 264800 03/06/09 02:50 Solvignon Spot++
  • 264797 03/06/09 02:46 Solvignon Harp scan results
  • 264794 03/06/09 01:38 Solvignon Target moved back to BeO+Optics
  • 264793 03/06/09 01:37 Solvignon NMR measurement: 63.17% (no beam)
  • 264792 03/06/09 01:32 Solvignon Target moved to He3 position for NMR measurement
  • 264789 03/06/09 00:01 R. Subedi Shift Summary (Swing)
  • Thursday

  • 264787 03/05/09 23:25 R. Subedi BB "magnet mismatch"
  • 264786 03/05/09 23:19 R. Subedi Beam status update
  • 264785 03/05/09 22:15 xiaohui Target moved to BeO
  • 264784 03/05/09 21:24 Yi Qiang Readings for Reference cell
  • 264783 03/05/09 20:24 R. Subedi Beam Status update
  • 264782 03/05/09 20:19 S. Sirca Reference cell gas/pressure
  • 264773 03/05/09 18:21 R. Subedi Beam off due to ARC 4 magnet trip
  • 264772 03/05/09 18:01 R. Subedi BeO 2x2.6 (MCC) raster
  • 264771 03/05/09 17:58 xiaohui NMR was done with beam off
  • 264770 03/05/09 17:57 xiaohui Target moved to 3He
  • 264768 03/05/09 17:53 xiaohui NMR screen shot
  • 264766 03/05/09 17:52 riordan Summary for March 4 Swing - March 5 Day
  • 264765 03/05/09 17:49 R. Subedi Harp IHA1H03A
  • 264764 03/05/09 17:47 R.Subedi Harp IHA1H03B
  • 264763 03/05/09 17:33 R. Subedi Beam Status
  • 264762 03/05/09 17:31 xiaohui Target moved to BeO
  • 264761 03/05/09 17:29 xiaohui NMR measurement
  • 264760 03/05/09 16:53 brads Reference cell had nitrogen @ 120psig
  • 264759 03/05/09 16:16 R. Subedi Beam in about 2 hourss
  • 264757 03/05/09 15:57 S. Sirca DAY shift summary
  • 264756 03/05/09 13:54 S. Sirca DOWN due to repeated arc magnet problems
  • 264752 03/05/09 13:50 S. Sirca spot++ with He3, 2uA, 5.5x6.5 raster + COMMENT
  • 264747 03/05/09 13:43 S. Sirca spot++ with He3, 2uA, 2.5x3.2 raster
  • 264743 03/05/09 13:37 S. Sirca spot++ with BeO, 2uA
  • 264740 03/05/09 12:49 S. Sirca Down due to arc magnet problems
  • 264739 03/05/09 12:41 Gao NMR measurement of 3He target polarization
  • 264728 03/05/09 12:07 ellie Sync Check Gaps (Run 2001)
  • 264727 03/05/09 12:05 mcnulty sync check breaks and DT = 21.5% for run 2001
  • 264724 03/05/09 11:59 S. Sirca Beam half-wave plate now OUT
  • 264721 03/05/09 11:46 Yi Qiang Please post the in beam reference cell pressure and gas type
  • 264716 03/05/09 11:12 ellie Golden Run changed to Runs 20546 (LHRS) and 1998 (BB)
  • 264715 03/05/09 11:10 S. Sirca Reference cell runs for kinematics G
  • 264710 03/05/09 11:02 ellie Sync Check Gaps (Run 1998)
  • 264705 03/05/09 10:16 gjin routine MWDC check
  • 264704 03/05/09 10:08 gjin MWDC hit efficiency
  • 264703 03/05/09 09:52 S. Sirca Hall A Tools after kinematics switch
  • 264702 03/05/09 09:48 ellie Sync Check Gaps (Run 1996)
  • 264699 03/05/09 09:33 S. Sirca Switch to kinematics G
  • 264694 03/05/09 08:47 ellie Sync Check Gaps (Run 1988)
  • 264693 03/05/09 08:46 mcnulty Breaks in sync check histos and 20% deadtime for run 1995
  • 264691 03/05/09 08:25 ellie Sync Check Gaps (Run 1994)
  • 264688 03/05/09 08:00 hyao Owl Shift Summary
  • 264686 03/05/09 07:27 hyao NMR=61.09%
  • 264682 03/05/09 07:15 Xinhu Yan Target Screen Shots
  • 264681 03/05/09 07:10 Xinhu Yan Absorption Spectrum
  • 264680 03/05/09 07:07 hyao halla tools
  • 264679 03/05/09 07:03 hyao spot++
  • 264660 03/05/09 03:54 Xinhu Yan Sync Check for run 1988
  • 264658 03/05/09 03:50 hyao Bigbite magnet tripped
  • 264652 03/05/09 03:10 Xinhu Yan Sync Check for run 1987
  • 264649 03/05/09 02:34 Xinhu Yan Sync Check for run 1986
  • 264644 03/05/09 00:40 hyao target NMR
  • 264639 03/05/09 00:11 Kalyan Discontinuity in sync check for run 1981
  • 264637 03/05/09 00:01 Kalyan Compton DAQ reboot
  • Wednesday

  • 264636 03/04/09 23:57 Gao Swing Shift Summary
  • 264633 03/04/09 23:22 Kalyan Re: BPM x and Y for run 20541
  • 264632 03/04/09 23:16 Kalyan BPM x and y for run 20541
  • 264631 03/04/09 23:11 Kalyan VDC efficiency
  • 264630 03/04/09 23:10 Kalyan Shift in Cerenkov ADCs.
  • 264623 03/04/09 21:59 Kalyan The BPM x and y looks narrower - run 20540
  • 264620 03/04/09 21:05 Kalyan Compton and parity DAQs started
  • 264612 03/04/09 20:44 A. Deur Raster size looks okay
  • 264610 03/04/09 20:41 A. Deur Beam is back
  • 264609 03/04/09 20:37 A. Deur NMR 62.83%
  • 264608 03/04/09 20:32 Yi Qiang Labview programs restarted
  • 264607 03/04/09 20:24 A. Deur NMR gui froze
  • 264606 03/04/09 20:23 A. Deur Laser 3 current
  • 264597 03/04/09 20:07 riordan Summary for Mar 3 Swing - Mar 4 Day
  • 264596 03/04/09 20:03 miham The Hall A tools
  • 264595 03/04/09 19:50 A. Deur Beam down since 5:42pm. No beam until at least 9h30
  • 264594 03/04/09 18:44 miham #20537 is a test run
  • 264592 03/04/09 18:35 Gao Beam spot check after beam is back
  • 264591 03/04/09 18:34 D. Parno LHRS golden run set to 20536
  • 264589 03/04/09 18:19 D. Parno Syncroc check -- Run 1977
  • 264588 03/04/09 18:14 A. Camsonne Edet Compton run 16200
  • 264587 03/04/09 17:58 D. Parno Syncroc check -- Run 1976
  • 264584 03/04/09 17:22 A. Camsonne Edet Compton run 16199
  • 264583 03/04/09 17:20 A. Camsonne Dark current Compton Edet
  • 264582 03/04/09 17:10 A. Camsonne Edet Compton run 16198
  • 264579 03/04/09 16:46 A. Deur Target NMR 62.68%
  • 264576 03/04/09 16:26 A. Deur Absorption spectrum
  • 264574 03/04/09 15:59 Saha Shift Summary
  • 264566 03/04/09 15:25 D. Parno Target moved to 3He
  • 264564 03/04/09 15:21 adaq spot++ result
  • 264560 03/04/09 15:17 brads All 3 aux DAQs restarted.
  • 264559 03/04/09 13:37 ellie Re: Left Arm Efficiencies (Run 20530)
  • 264558 03/04/09 13:32 posik BB cer 1917-1919 analysis
  • 264557 03/04/09 13:19 O. Glamazdin Hall A tools strip chart
  • 264556 03/04/09 13:07 R. Michaels absorption spectrum
  • 264555 03/04/09 12:40 R. Michaels target to BeO
  • 264546 03/04/09 12:10 R. Michaels NMR (62.54)
  • 264545 03/04/09 12:05 R. Michaels adaql3 reboot, revived charts
  • 264544 03/04/09 11:21 ellie Left Arm Efficiencies (Run 20530)
  • 264543 03/04/09 11:04 gjin routine MWDC check (hit efficiency)
  • 264542 03/04/09 10:52 gjin routine MWDC check
  • 264541 03/04/09 10:31 ellie Sync Check1(straight lines) Problem for Run 1965
  • 264540 03/04/09 10:04 ellie PID Detectors (Good Electrons) for posF90 kinematics
  • 264535 03/04/09 09:09 ellie Golden Runs set to 20530 (LHRS) and 1972 (BB)
  • 264534 03/04/09 08:38 riordan Auxiliary DAQs Started
  • 264529 03/04/09 08:26 R. Michaels dipole settled, spot++, no N2
  • 264528 03/04/09 08:11 R. Michaels near-term plan
  • 264527 03/04/09 08:04 Gao Owl Shift Summary
  • 264525 03/04/09 08:02 R. Michaels L-HRS polarity change
  • 264524 03/04/09 07:42 Ewing/Fansle
  • 264523 03/04/09 07:15 Nelyubin DBB lost syncroc at run #1971
  • 264522 03/04/09 07:00 riordan Going to Controlled Access
  • 264515 03/04/09 06:09 Nelyubin DBB syncroc is better
  • 264510 03/04/09 05:33 Nelyubin Lost syncroc DBB
  • 264491 03/04/09 03:08 cusanno Spot++ (run 20523)
  • 264489 03/04/09 03:06 cusanno Absorbtion spectrum
  • 264488 03/04/09 03:03 cusanno NMR (62.41)
  • 264484 03/04/09 02:33 cusanno Spot++ (run 20523)
  • 264480 03/04/09 01:47 Nelyubin Target polarization and the holding field Graph
  • 264479 03/04/09 01:43 Nelyubin Strip Charts Target Oven and Cell temperature
  • 264478 03/04/09 01:42 xqian sync problem
  • 264473 03/04/09 01:26 Nelyubin Absorption Spectrum
  • 264448 03/04/09 00:09 jpchen Swing shift summary
  • Tuesday

  • 264444 03/03/09 23:53 xiaohui Target charts
  • 264441 03/03/09 23:50 Xiaohui Target moved to N2 120psig
  • 264434 03/03/09 23:18 Chiranjib L1A vs clock check
  • 264431 03/03/09 23:15 Chiranjb Dead time check
  • 264428 03/03/09 22:53 Bo Zhao L-arm VDC efficiency
  • 264427 03/03/09 22:46 jpchen BB deadtime reasonable (17%)
  • 264424 03/03/09 22:42 jpchen Completed Kin. F, starting kin. D
  • 264423 03/03/09 22:41 Xiaohui NMR measurement before production data taking
  • 264422 03/03/09 22:36 Xiaohui Target moved to 3He
  • 264419 03/03/09 22:34 Xiaohui N2 pressure in beam is 122psig
  • 264418 03/03/09 22:31 jpchen N2 run for kin F.
  • 264412 03/03/09 22:21 jpchen Left VDC HVs reset
  • 264409 03/03/09 22:11 jpchen BB Hvs are on
  • 264408 03/03/09 22:03 xiaohui Target moved to N2 reference cell. 120psig
  • 264407 03/03/09 22:00 jpchen working on turning on HVs for BB chambers and LHRS VDCs
  • 264406 03/03/09 21:59 jpchen raster check
  • 264404 03/03/09 21:44 jpchen Moved to BeO first befofe N2, checking raster
  • 264403 03/03/09 21:44 xiaohui Target moved to BeO to check raster size
  • 264402 03/03/09 21:42 O. Glamazdin Moller measurements result
  • 264401 03/03/09 21:42 jpchen Moller completed, beam tuned for production. Moveing to N2
  • 264400 03/03/09 21:41 xiaohui Target moved to reference cell
  • 264399 03/03/09 21:36 jpchen new structure should be protons
  • 264398 03/03/09 21:33 jpchen LHRS trigger rates + vs -.
  • 264397 03/03/09 21:27 riordan Summary for March 2 Swing - March 3 Day
  • 264396 03/03/09 21:04 jpchen scaler BB T2 rates on Sunday and today are about the same
  • 264395 03/03/09 20:45 jpchen raster for run 20502
  • 264386 03/03/09 19:15 A. Camsonne Compton Edet test long run 16190 for efficiency
  • 264385 03/03/09 19:12 A. Camsonne Compton Electron detector dark current
  • 264384 03/03/09 19:00 brads Happex (and Compton FADC) runs restarted
  • 264383 03/03/09 18:54 A. Camsonne Compton Edet test run 16189
  • 264382 03/03/09 18:48 A. Camsonne Compton Edet test run 16188
  • 264381 03/03/09 18:46 brads Re-enabled Left Q3 in alarm handler
  • 264380 03/03/09 18:46 A. Camsonne Run Edet test 16187
  • 264379 03/03/09 18:41 A. Camsonne Compton electron detector dark current
  • 264378 03/03/09 18:37 brads ROC5 sync issue fixed (needed power cycle)
  • 264377 03/03/09 18:33 jpchen HVs for LHRS VDC tripped and will stay off until Moller over
  • 264372 03/03/09 18:23 Jack RICH gas supply changed to UHP grade Nitrogen
  • 264371 03/03/09 18:13 Parno/Camson Compton asymmetry measurement
  • 264370 03/03/09 18:05 jpchen Stopped data taking for MOLLER, moved to EMPTY, BB CH HV off
  • 264369 03/03/09 18:00 xiaohui Target moved to empty
  • 264368 03/03/09 17:57 D. Parno Sync check for Run 1949
  • 264365 03/03/09 17:39 brads Run extended to Mar. 16
  • 264364 03/03/09 17:36 D. Parno ROCs and TSes rebooted
  • 264363 03/03/09 17:33 brads Linac energies around Feb 28
  • 264359 03/03/09 17:00 D. Parno LHRS golden run set to 20499
  • 264358 03/03/09 16:55 D. Parno LHRS VDC Efficiency Capture
  • 264356 03/03/09 16:45 riordan Deadtime Back to Reasonable
  • 264354 03/03/09 16:44 D. Parno HAPPEX/Compton restart
  • 264353 03/03/09 16:43 A. Camsonne Compton EPICS slow control rebooted
  • 264344 03/03/09 16:04 X. Zheng Shift Summary (Day)
  • 264343 03/03/09 16:03 X. Zheng sync problem remained for 1952
  • 264342 03/03/09 15:58 D. Parno High deadtime for Run 1952
  • 264341 03/03/09 15:58 X. Zheng Run list for positron Kinematics H
  • 264340 03/03/09 15:50 X. Zheng Rebooted ROCs 5, 8, 9, 10 etc, and changed LHRS p0 to 1.13 (#F)
  • 264339 03/03/09 15:49 riordan BigBite ROCS and TS rebooted
  • 264339 03/03/09 15:49 D. Parno High offline BB deadtime
  • 264338 03/03/09 15:44 Moffit Re: syncroc5
  • 264335 03/03/09 15:35 D. Parno Syncroc 5 doesn't increment in Run 1950
  • 264334 03/03/09 15:29 brads re: ROC Sync Bad (probably OK)
  • 264331 03/03/09 15:14 riordan ROC Sync Bad
  • 264330 03/03/09 15:13 X. Zheng BB sync check not right since run 1948
  • 264328 03/03/09 15:02 D. Parno No change in Compton asymmetry after 2/18 half waveplate switch
  • 264325 03/03/09 14:45 D. Parno Compton /HAPPEX restart
  • 264324 03/03/09 14:29 Jack (Techs) Left Q3 current instability
  • 264319 03/03/09 13:43 D. Parno Compton rates are OK
  • 264318 03/03/09 13:33 X. Zheng spot++ on polarized 3He
  • 264317 03/03/09 13:32 D. Parno New structure in LHRS S2 ADCs
  • 264310 03/03/09 12:50 golge Input on LHRS # 20494 BB #1948
  • 264309 03/03/09 12:50 X. Zheng spot++ for N2 run
  • 264306 03/03/09 12:47 Xinhu Yan NMR Frequency Sweep Diagram
  • 264305 03/03/09 12:34 X. Zheng harp scan 160x157um
  • 264302 03/03/09 12:31 X. Zheng spot++ on BeO (2x2mm)
  • 264291 03/03/09 12:09 riordan BigBite Thresholds Set
  • 264290 03/03/09 12:08 X. Zheng Tuned beam
  • 264285 03/03/09 11:08 gjin routine MWDC check
  • 264284 03/03/09 10:53 brads Switched to full C4F8O tank on BB Cerenkov
  • 264281 03/03/09 10:00 X. Zheng Hall is going to controlled access
  • 264280 03/03/09 09:40 Xinhu Yan Target moved from 3He to Empty target
  • 264279 03/03/09 09:07 riordan Need to go to controlled access for Left Q3 work
  • 264276 03/03/09 08:05 Lindgren Owl Shift Summary
  • 264274 03/03/09 08:01 riordan Q3 Current Not Improved
  • 264273 03/03/09 06:55 Nelyubin Absorption Spectrum
  • 264272 03/03/09 06:51 Nelyubin Target Holding Field and Polarization Status
  • 264271 03/03/09 06:48 Nelyubin Target Oven and Cell Temperature
  • 264270 03/03/09 06:10 lindgren left arm high voltage
  • 264268 03/03/09 05:43 Lindgren No Beam
  • 264257 03/03/09 03:48 Lindgren still no beam.
  • 264256 03/03/09 03:04 Lindgren He and N2 levels
  • 264255 03/03/09 03:02 Lindgren Hall A Tools
  • 264252 03/03/09 00:42 Lindgren Owl shift starting
  • 264251 03/03/09 00:26 Yi Qiang Previous reference cell pressure was 113 psig
  • 264250 03/03/09 00:18 Nelyubin reference cell filled N2 at 120 psig
  • Monday

  • 264248 03/02/09 23:14 Xiaohui NMR measurement for swing shift
  • 264247 03/02/09 22:50 brads Left Q3 alarm DISABLED in alarm handler
  • 264246 03/02/09 22:40 brads BigBite DAQ fixed
  • 264237 03/02/09 21:47 riordan Summary for March 1 Swing - March 2 Day
  • 264224 03/02/09 21:00 riordan Swing Shift on Standby
  • 264219 03/02/09 20:43 riordan Left Q3 current unstable
  • 264218 03/02/09 20:31 A. Camsonne Compton electron detector dark current
  • 264209 03/02/09 19:40 brads BigBite set to 710A
  • 264208 03/02/09 19:27 Jack (Techs) BigBite power supply water leak
  • 264207 03/02/09 19:20 Jack (Techs) Left VDC discriminator threshold
  • 264206 03/02/09 18:19 xiaohui BigBite magnet Off
  • 264205 03/02/09 18:18 brads BigBite ramped down
  • 264204 03/02/09 18:16 brads LHRS & RHRS VDC HVs restored.
  • 264201 03/02/09 16:53 Parno / Cams Compton: Polarization drop over weekend
  • 264200 03/02/09 16:42 brads LHRS set to 1.270 GeV/c
  • 264199 03/02/09 16:17 Yi Qiang Routine Laser diode temperature check
  • 264198 03/02/09 15:59 Xinhu Yan Day Shift Summary
  • 264194 03/02/09 14:40 Xinhu Yan Target NMR Frequency Sweep Control Screen
  • 264193 03/02/09 14:37 Xinhu Yan Absorption spectrum
  • 264192 03/02/09 14:21 Xinhu Yan Hall A Tool Screen Shot
  • 264191 03/02/09 13:53 folts Left Dipole polarity
  • 264190 03/02/09 13:46 brads No beam until ~9pm
  • 264189 03/02/09 13:04 brads re: iocha11 and iocha22 (comm restored)
  • 264188 03/02/09 12:42 Jin/Yi Target Masing Effect Test
  • 264177 03/02/09 11:38 brads Missing EDTM in BigBite triggers (solved)
  • 264176 03/02/09 11:18 Kalyan Re: BCM Calibration: BCM1 and 2 vs 0L02 -- new plot
  • 264175 03/02/09 10:21 gjin routine MWDC check
  • 264174 03/02/09 08:49 brads Ramped HRS to 0 for polarity change
  • 264173 03/02/09 08:48 brads (Tech) Left VDC discrim. level drifting
  • 264169 03/02/09 04:54 cusanno Laser Absorption Spectrum
  • 264168 03/02/09 04:52 cusanno NMR (62.92)
  • 264167 03/02/09 04:19 cusanno Hall A tools
  • 264156 03/02/09 01:51 R. Lindgren Summary
  • 264155 03/02/09 00:32 brads re: ioc11 and ioc22
  • 264154 03/02/09 00:28 Lindgren Owl shift start
  • 264153 03/02/09 00:18 brads re: ioc11 and ioc22
  • 264152 03/02/09 00:01 bogdanw shift summary
  • Sunday

  • 264150 03/01/09 23:58 bogdanw ioc11 and ioc22
  • 264148 03/01/09 23:38 Jin Huang JLab laser over temperature; check output; power on again
  • 264146 03/01/09 23:30 brads HRS HV communication lost
  • 264145 03/01/09 23:18 A. Rakhman Hall A Status
  • 264144 03/01/09 23:15 brads CHL glitch -- beam down for several hours
  • 264141 03/01/09 22:23 A. Rakhman Target status
  • 264140 03/01/09 22:18 A. Rakhman target
  • 264139 03/01/09 22:18 A. Rakhman all 3 lasers have been switched off
  • 264138 03/01/09 22:16 A. Rakhman warning massege re pumping chamber temp, check oven and laser
  • 264137 03/01/09 22:03 A. Rakhman Beam Status
  • 264134 03/01/09 21:17 A. Rakhman NMR sweep swing shift 03/01/09 20:22 ( 59.87 % transv -)
  • 264133 03/01/09 21:11 A. Rakhman NMR sweep swing shift 03/01/09 20:22 ( 62.24% transv -)
  • 264132 03/01/09 21:03 A. Rakhman Target screen snapshot
  • 264131 03/01/09 21:02 A. Rakhman Laser Absorption Spectrum
  • 264130 03/01/09 21:01 ycwang RUN 1928 and 20480 are cosmic run, almost no beam for these runs
  • 264125 03/01/09 20:37 Jin Huang Spin Flip to 90 Degree/ Transverse+
  • 264122 03/01/09 20:20 ycwang Run 1927 and 20479 stopped for Spin Flip
  • 264107 03/01/09 18:38 posik Happex and comptons restarted
  • 264104 03/01/09 18:02 A. Rakhman Target screen snapshot
  • 264103 03/01/09 18:02 ycwang run 1925 and 20477 stopped
  • 264102 03/01/09 18:02 posik kin G270 golden run->20476
  • 264101 03/01/09 18:01 A. Rakhman Laser Absorption Spectrum 03_01_2009
  • 264100 03/01/09 18:00 A. Rakhman 1st_NMR_03_01_Swing_shift ( 63.4 % )
  • 264097 03/01/09 17:46 riordan Summary for Feb 28 Swing - March 1 Day
  • 264096 03/01/09 17:31 A. Rakhman Target NMR measurement
  • 264095 03/01/09 17:30 ycwang RUN Coordinator reboot adaql2. Now we are taking data.
  • 264093 03/01/09 17:22 riordan Summary for Feb 28 Swing - March 1 Day
  • 264091 03/01/09 16:49 ycwang adaql2 does not work
  • 264090 03/01/09 16:34 D. Parno HAPPEX/Compton Restart
  • 264088 03/01/09 16:03 cusanno kinematics G production started
  • 264086 03/01/09 16:03 cusanno Day Shift Summary
  • 264084 03/01/09 16:01 garibaldi He-3 target in
  • 264079 03/01/09 15:14 garibaldi N2 reference cell in
  • 264078 03/01/09 15:13 cusanno Kinematics G
  • 264077 03/01/09 15:13 gjin routine MWDC check
  • 264071 03/01/09 14:25 D. Parno Compton /HAPPEX Restart
  • 264070 03/01/09 14:13 cusanno beam back
  • 264069 03/01/09 14:13 cusanno BB VDC trip
  • 264068 03/01/09 14:10 cusanno Hall A Tools and BCM
  • 264065 03/01/09 13:35 cusanno Beam off
  • 264062 03/01/09 13:07 garibald Target Laser absorption spectrum
  • 264059 03/01/09 12:58 F. Garibaldi 1st_NMR_02_01_day_shift
  • 264056 03/01/09 12:41 posik Happex and both comptons restarted
  • 264055 03/01/09 12:38 A. Camsonne Compton ROC not booted
  • 264044 03/01/09 12:09 posik BB cer ADC shift
  • 264041 03/01/09 10:45 posik HAPPEX and FADC compton restarted
  • 264038 03/01/09 10:36 cusanno Beam is back
  • 264037 03/01/09 10:04 cusanno Beam line problems
  • 264036 03/01/09 09:56 brads (Tech) LHRS Q1 LHe level dropping to yellow
  • 264033 03/01/09 09:39 cusanno beam quality for the last 90 min
  • 264032 03/01/09 09:35 cusanno BB HV trip, channel 14.6
  • 264031 03/01/09 09:29 cusanno beam back
  • 264030 03/01/09 09:27 cusanno Beam trip
  • 264029 03/01/09 09:14 brads Run 1916 is probably junk
  • 264027 03/01/09 09:04 cusanno Left Q1 recovered
  • 264025 03/01/09 08:59 brads HV crate turned off, now back on
  • 264022 03/01/09 08:47 posik no compton rate
  • 264021 03/01/09 08:40 cusanno BB HV Alarm
  • 264020 03/01/09 08:39 cusanno Controlled Access
  • 264017 03/01/09 08:29 cusanno Left Q1 trip
  • 264013 03/01/09 08:02 Fatiha Shift Summary
  • 264012 03/01/09 07:46 Fatiha Beam current for the last 2h
  • 264011 03/01/09 07:45 R. Michaels NMR (63.87)
  • 264006 03/01/09 06:57 Kalyan BB re-timing is good : opitmal trigger condition
  • 264005 03/01/09 06:56 brads BB golden rootfile -> run 1914
  • 264004 03/01/09 06:23 brads New FADC and Happex runs started.
  • 264003 03/01/09 06:22 brads (late entry) Back to beam permit, MCC struggling with vacuum problems
  • 264002 03/01/09 06:21 Fatiha Back to production
  • 264001 03/01/09 06:20 Kalyan Shower threshold vs Shower X
  • 263998 03/01/09 06:17 brads re: timing of PSh discriminator in the T6 trigger
  • 263997 03/01/09 06:07 brads T2 timing shift reverted (fixed the timing glitch seen in runs 1911, 1912)
  • 263996 03/01/09 06:03 Kalyan preshower_E vs shower_E after trigger modification
  • 263995 03/01/09 05:58 Fatiha Hall in controlled access
  • 263994 03/01/09 05:47 Fatiha Beam quality for the last 2h
  • 263991 03/01/09 05:28 R. Michaels Absorption Spectrum
  • 263988 03/01/09 05:23 R. Michaels gas C TDC plots
  • 263981 03/01/09 04:36 brads T1 is plugged back into the TS (but should normally be prescaled away)
  • 263980 03/01/09 04:34 Fatiha spot++ for production running
  • 263979 03/01/09 04:34 brads Added PSh discriminator to the T6 trigger
  • 263978 03/01/09 04:31 brads New Compton FADC and Happex runs started
  • 263975 03/01/09 04:23 lindgren tuning raster
  • 263974 03/01/09 04:20 brads Adjusted T2 arrival time at the trigger supervisor
  • 263973 03/01/09 04:19 lindgren tuning raster size
  • 263972 03/01/09 04:12 Fatiha 2*2 spot size with BeO
  • 263971 03/01/09 04:11 R. Michaels move to He3
  • 263960 03/01/09 04:04 Fatiha Harp Scan results
  • 263959 03/01/09 04:00 R. Michaels target to Optics
  • 263958 03/01/09 03:57 Kalyan Run 1910 is not a production
  • 263957 03/01/09 03:57 brads Run 1910 was just a test run
  • 263956 03/01/09 03:56 brads/kalyan Controlled Access (opportunistic)
  • 263954 03/01/09 03:39 R. Michaels target to empty
  • 263953 03/01/09 03:10 Fatiha Update on the beam situation
  • 263952 03/01/09 02:38 R. Michaels NMR (63.46)
  • 263950 03/01/09 01:25 R. Michaels more on Qasy (it looks ok)
  • 263949 03/01/09 01:13 Kalyan Re:Cerenkov ADC spectrum shift
  • 263948 03/01/09 01:05 Fatiha Beam down since 00:42
  • 263945 03/01/09 00:19 posik cer results for runs 1848,1849,1851
  • 263943 03/01/09 00:02 bogdanw shift summary